Suppose Bob owns 2,000 shares of Company X and 10,000 shares of Company Y. The total value of Bob's holdings of these two companies is $340,000.
Suppose Frank owns 8,000 shares of Company X and 6,000 shares of Company Y. The total value of Franks holdings of these two companies is $340,000.
a. Write equations for Bob and Frank's holdings. Use the variables X and Y to represent the values of shares of Company X and Company Y.
b. Solve for the value of a share of Company X and Company Y. Show your work so you can get partial credit even if your final answer is wrong.
I do not understand how to get "b": solving the value of a share compnay?

Answer by stanbon(48575)

Suppose Bob owns 2,000 shares of Company X and 10,000 shares of Company Y. The total value of Bob's holdings of these two companies is $340,000.
Eq #1: 2000x + 10000y = 340000
---------------------------

Suppose Frank owns 8,000 shares of Company X and 6,000 shares of Company Y. The total value of Franks holdings of these two companies is $340,000.
Eq #2: 8000x + 6000y = 340000
---------------------------
a. Write equations for Bob and Frank's holdings. Use the variables X and Y to represent the values of shares of Company X and Company Y.
b. Solve for the value of a share of Company X and Company Y. Show your work so you can get partial credit even if your final answer is wrong.
----
Solve by any method you know to get:
Eq #1: 2000x + 10000y = 340000
Eq #2: 8000x + 6000y = 340000
---
Divide thru both of those equations by 1000 to get:
Eq #1: 2x + 10y = 340
Eq #2: 8x + 6y = 340
-------------------------
Multiply thru the 1st equation by 4:
Eq #1: 8x + 40y = 1360
Eq #2: 8x + 6y = 340
-------------------------
Subtract the 2nd Eq from the 1st to solve for "x":
34y = 1020
y = 30
-------------
Substitute into 2x + 10y = 340 to solve for "x":
2x + 10*30 = 340
2x = 40
x = 20
=======================================
Final Answer:
X = $20
Y = $30
====================
Cheers,
Stan H.